pooja195 (pooja195):

Solve the system using linear combinations. 3m + 2n = 10 4m – n = –1 I need someone to tell me if im doing this right im not sure if I get this or not =/

pooja195 (pooja195):

ok so I should start by multiplying right?

Nnesha (nnesha):

yep \(\rm\color{green}{o^\wedge\_^\wedge o}\)

pooja195 (pooja195):

\[-4(3m+2n=10)\] \[3(4m – n = –1)\]

OpenStudy (freckles):

that works but I think you took the harder route

OpenStudy (solomonzelman):

you can do matrix 9just set it up and solve), substitution (rearrange the second equation), or elimination (multiply the equation to make the values of one of variables be additive inverses in the 2 equations)

pooja195 (pooja195):

=/ im confused xD and i realize i dont get this :(

OpenStudy (freckles):

you could did what you did... (or what I thought was easier which was to multiply the second equation by 2 then add the equations together ) But your way works as well. You just have to multiply both equations be a number which is not too much extra work.

OpenStudy (freckles):

distribute your -4 to your first equation and your 3 to your second then add your equations

OpenStudy (solomonzelman):

\(\large\color{slate}{ 3m + 2n = 10 }\) \(\large\color{slate}{ 4m - n = -1 }\) >>> \(\large\color{slate}{ 4m +1 = n }\) \(\large\color{slate}{ 3m + 2(4m +1 ) = 10 }\)

OpenStudy (solomonzelman):

not so hard:P

OpenStudy (solomonzelman):

then solve for n as you find M

OpenStudy (solomonzelman):

I see a sub as the quickest approach

Nnesha (nnesha):

you have to multiply by something so you can cancel out one variable first equation have positive 2n so you can multiply 2nd equation by just 2 \(\rm\color{green}{o^\wedge\_^\wedge o}\) \(\rm\color{green}{o^\wedge\_^\wedge o}\)

pooja195 (pooja195):

\[3m+8m+2=10\]

pooja195 (pooja195):

and then 11m+2=10 11m=7??????

OpenStudy (solomonzelman):

10-2 is 8 (not 7) yes, apparently m=11/8

OpenStudy (solomonzelman):

I mean m=8/11

pooja195 (pooja195):

oops sorry right and then now i would just plug that into one of the equations right?

OpenStudy (solomonzelman):

yes, I would plug m=8/11 into both of the equations to also verify it

OpenStudy (solomonzelman):

https://www.desmos.com/calculator/pvyxhl5bzg verification

pooja195 (pooja195):

Oh ok thanks i got it :D

OpenStudy (solomonzelman):

:)